Let f(x) = e x − 3 define a real-valued function. Using an initial guess of w0 = 1, perform one iteration of Newton’s method to approximate the zero of f. Compute and simplify the error of your approximation.
Answer #1

G1iven funct ion ts f)=e-3 Condition iis ,=1 The initia! The newtons teration foremula is e 3 (e-3) nt 1 e 3 e e +3 nt1 +3 F1.1036 I036 11036 XO.1036 t 3 F1.0986 1.1036 e fon nE2 3 +3 1-0986 (oe6-1) 0986 +3 1.0986 .e0.0186) +3 = 1098 60 0986 an apptThe erercore = Exaet ratuf approximation ralmt 1.0986122886 811-10986G 70 -5 1228866811 ID we calcul ate appreoximate ret CoP